OpenStudy (anonymous):

To test that the solution to the equation below is x = 4, which of the following calculations should be performed?

OpenStudy (elonasushchik):

Hminus 3 to both sides is next. - 2/4x + 3 = 1 - 2/4x = - 2 - 1/2x = - 2 (2/1)1/2x = (2/1)2 x = 4.

OpenStudy (anonymous):

thanks

OpenStudy (elonasushchik):

Medal?!